# RealCamo: Boosting Real Camouflage Synthesis with Layout Controls and Textual-Visual Guidance

### v1

@@ 3.4 Texture-Oriented Background Retrieval
Compared to generic image synthesis, camouflaged image generation requires backgrounds that are quite complex and sometimes unconventional. Although LAKE-RED (CVPR2024LAKERED) retrieves background embeddings from a pretrained codebook, such representations learned from generic image distributions lack camouflage-specific priors. We therefore introduce a texture-oriented background retrieval (TOBR) module that uses camouflaged images as a domain-specific background knowledge base. As illustrated in Algorithm 1, given a target foreground object, a CLIP image encoder is adopted to extract texture-aware features for the object and all candidates in the knowledge base, with the top-$k$ most similar images in the embedding space are retrieved and then incorporated into multi-modal prompt construction to guide the camouflage generation process.
@@ 3.5 Textual-Visual Condition Generator
Textual conditioning is central to T2I generation. Although CVPR2025CamoAny enhances textual prompts via prompt engineering, the incorporated background semantics remain limited. While detailed scene descriptions have been shown to improve the realism of camouflaged image synthesis (AAAI2026CTCIG), acquiring scene-specific fine-grained descriptions is often labor-intensive. We therefore propose a textual–visual condition generator (TVCG) that constructs a multi-modal prompt by combining a unified fine-grained task description with high-quality background visual embeddings. This design offers richer guidance for generation and improves visual similarity between the foreground target and the background. The multi-modal prompt comprises three components: (i) fine-grained textual task description, (ii) object category information, and (iii) reference background visual cues.
Fine-Grained Textual Task Description. Instead of using overly simplistic prompts (e.g., “image of a fish Camouflaged with Background” (CVPR2025CamoAny)) or labor-intensive, scene-specific descriptions (e.g., “A fish with mottled green and brown patterns and elongated appendages in a coral reef ecosystem.” (AAAI2026CTCIG)), we adopt a unified fine-grained textual task description “A realistic image of an object blending into its surroundings, where the background shares similar colors, textures, and patterns with the object, making it hard to distinguish. Natural lighting, photorealistic, seamless camouflage, high detail.” as the textual guidance in generation. This unified description improves generation quality while eliminating the need for manually crafting scene-specific prompts, and is encoded into textual tokens $c_{tex}$ via a CLIP text encoder.
Object Category Information. For the given image, we adopt the [CLS] token processed by a CLIP image encoder as a compact category representation of the foreground object, avoiding explicit class identification while preserving informative semantic guidance.
@@
Reference Background Visual Cues. To provide richer background guidance for the camouflage generation process, we incorporate the samples retrieved by TOBR as visual reference knowledge into the multi-modal prompt. Specifically, we employ a CLIP image encoder to extract visual embeddings from the input image $\mathrm{I}$ and its top-$k$ TOBR-retrieved samples $\{S_{j}\}_{j=1}^{k}$. Guided by the foreground mask $\mathrm{M}$, the foreground region of the input image is preserved, while the background regions are fused with the retrieved visual knowledge to construct background-aware visual cues $c_{vis}$. This process can be formulated as:
@@
$c_{vis}=\left\{\begin{aligned} &e_{t}\cdot\mathrm{M}+\frac{1}{k+1}\cdot(e_{t}+\sum_{j=1}^{k}e_{j})\cdot(\mathbf{1}-\mathrm{M}),\ if\ training\\ &e_{t}\cdot\mathrm{M}+\frac{1}{k}\cdot\sum_{j=1}^{k}e_{j}\cdot(\mathbf{1}-\mathrm{M}),\ if\ inference.\end{aligned}\right$. (5)
The final multi-modal conditional prompt is the concatenation of $_{txt}, cc$$_{cls}$, and $c_{vis}$ along the sequence dimension.
Figure 4: Qualitative comparison with existing SOTA methods including style transfer based (column 3 to 4), out-painting based (column 5 to 9), and text-driven (column 10) approaches. Our results are demonstrated in the rightmost column with red box.
Table 1: Quantitative comparison with SOTA methods on LAKE-RED datasets. The 1st, 2nd and 3rd best results are highlighted in red, green and blue. $\uparrow$ / $\downarrow$ represents the higher / lower the better. ST: Style transfer based. OP: Out-paining based. TD: Text-driven.
@@ 3.6 Training Objective
In RealCamo, the Stable Diffusion (SD) and the CLIP text and image encoders are all frozen, leaving ControlNet as the only trainable component. The training objective follows Equation 2, where the original control $c^{f}$ is replaced by layout controls $\{c_{\mathrm{cst}},c_{\mathrm{dep}},c_{\mathrm{hed}}\}$, and the textual condition $c^{t}$ is replaced by multi-modal conditions $\{c_{\mathrm{txt}},c_{\mathrm{cls}},c_{\mathrm{vis}}\}$. This design enables effective controllable camouflage synthesis while preserving the generalization of the pretrained SD.
@@ 3.7 Background-Foreground Distribution Divergence
To more effectively quantify the degree of camouflage in synthesized images, we propose a background–foreground distribution divergence metric, denoted as $\mathrm{KL}_{BF}$, based on KL divergence. It measures visual similarity between the generated background and the foreground target via their pixel-value distributions, where lower values indicate stronger camouflage effectiveness. Specifically, given an image $\mathrm{I}$ and its foreground mask $\mathrm{M}$, we compute the KL divergence between background and foreground distributions for each RGB channel and average them to obtain $\mathrm{KL}_{BF}$. Formally, the $\mathrm{KL}_{BF}$ can be calculated as:
$\mathrm{KL}_{c}$ $=KL(\mathrm{I}_{c}\cdot(\mathbf{1}-\mathrm{M}),\mathrm{I}_{c}\cdot\mathrm{M}),\ c\in\{R,G,B\},$ (6)
$\mathrm{KL}_{BF}$ $=Avg(\mathrm{KL}_{R},\mathrm{KL}_{G},\mathrm{KL}_{B}).$ (7)
Figure 5: Qualitative results of ablation study on main components of our proposed RealCamo framework.
Table 2: Quantitative results of ablation study on main components. The best results are highlighted in bold. $\uparrow$ / $\downarrow$ represents the higher / lower the better. For layout controls, $\times$ means only a foreground target with white background is used, without depth and HED controls. For textual-visual condition, $\times$ means only simple textual prompt is used, without additional visual tokens.
@@ 4.1 Experimental Settings
Datasets and Metrics. We conduct experiments on the LAKE-RED dataset (CVPR2024LAKERED), which contains 4,040 training images and 19,419 images for evaluation, including 6,473 camouflaged, salient, and general images, respectively. Following CVPR2024LAKERED, we adopt FID (ICLR2018FID) and KID (NIPSl2017KID) to assess the overall quality of generated images. In addition, SSIM (TIP2004SSIM) is employed to measure structural consistency between generated images and their original references, while our proposed $\mathrm{KL}_{BF}$ metric is used to quantify the effectiveness of visual camouflage.
Implementation Details. The proposed framework is built upon Stable Diffusion 1.5 (CVPR2022LDM) augmented with ControlNet (ICCV2023ControlNet). During training, all input images, control conditions, and corresponding masks are resized to $512\times 512$, and a batch size of 4 is used. We set the control scale to 1.0 and train the model for 600,000 iterations with a learning rate of 1e-5. The generated images are also in a resolution of $512\times 512$ and are resized to their original size for evaluation. All experiments are conducted on a single NVIDIA L20 GPU.
@@ 4.2 Comparison with the State-of-the-arts (SOTAs)
We compare our RealCamo with 12 SOTAs, including five style transfer based methods (AB (IPOL2016AB), CI (TOG2010Chu), AdaIN (ICCV2017AdaIN), DCI (AAAI2020DCI), and LCGNet (TMM2022LCGNet)), six out-painting based methods (TFill (CVPR2022TFill), LDM (CVPR2022LDM), RePaint-L (CVPR2022Repaint), LAKE-RED (CVPR2024LAKERED), FACIG (ICME2025FACIG), and CamoAny (CVPR2025CamoAny)), and one recently text-driven method (CT-CIG (AAAI2026CTCIG)). We re-evaluate the benchmark results provided by LAKE-RED and CT-CIG. Due to a lack of available resources, we reproduce CamoAny with the same settings as ours, and directly copy the results from the original paper for FACIG.
Table 5: Performance of MVANet (CVPR2024MVANet) on four COD benchmark datasets with different training data. The best results are highlighted in bold. $\uparrow$ / $\downarrow$ represents the higher / lower the better. Original: The traditional COD training data. Gen SOD&GOD: Synthesis images by RealCamo of the salient and general subsets in LAKE-RED evaluation set. Gen COD: Synthesis images by RealCamo of the LAKE-RED training set (i.e., our constructed SynCOD12K dataset).
@@
Gen SOD&GOD 0.905 0.927 0.846 0.796 0.031 0.826 0.842 0.760 0.717 0.079 0.865 0.896 0.784 0.719 0.035 0.879 0.897 0.821 0.769 0.048
Gen COD 0.924 0.949 0.889 0.871 0.019 0.882 0.916 0.856 0.826 0.049 0.896 0.934 0.850 0.818 0.021 0.891 0.919 0.860 0.824 0.037
Qualitative Comparison. Fig. 4 shows qualitative comparison between our proposed RealCamo and other methods on camouflaged, salient, and generic objects. The style transfer based LCGNet embeds targets into given backgrounds, achieving strong visual camouflage but severely altering target semantics. In contrast, out-painting based methods often produce insufficient camouflage or cluttered, semantically implausible backgrounds, degrading realism. Although recent text-driven CT-CIG generates visually convincing results, they fail to preserve original targets and thus require re-annotation, limiting their usefulness for downstream COD tasks. In comparison, our RealCamo achieves both high visual similarity and strong foreground–background semantic consistency while preserving target integrity, resulting in more realistic and practically valuable camouflaged images.
Quantitative Comparison. Quantitative results are summarized in Tab. 1. Compared to early style transfer based methods, latest out-painting based and text-driven approaches achieve obviously better image quality, indicated by lower FID and KID. Owing to the introduction of explicit layout controls and textual–visual multi-modal conditioning, our RealCamo attains highly competitive generation performance. Although it does not achieve the lowest $\mathrm{KL}_{BF}$ score, our method balances visual camouflage with structural and semantic fidelity, producing more realistic images. By revisiting Fig. 4, methods with lower $\mathrm{KL}_{BF}$ values (e.g., CamoAny and TFill) indeed yield stronger visual camouflage, further validating our proposed $\mathrm{KL}_{BF}$ as an effective metric for quantifying camouflage effectiveness.
User Study. We further conduct a user study to comprehensively evaluate the quality of synthesized camouflaged images based on human feedback. Detailed experimental settings, evaluation criteria, and statistical results and analyses will be reported in Appendix C.
@@ 4.3 Ablation Study
We conduct ablation studies to validate the proposed components. Due to computational constraints, all variants are trained for 100,000 iterations, and evaluation is performed on 6,000 images uniformly sampled from three test subsets. Results in Tab. 2 and Fig. 5 show that explicit layout controls improve structural consistency and overall image quality, supporting our claim that “the semantics of images are largely determined by their structural layout”. Incorporating textual–visual condition further enhances visual realism, as well as foreground–background visual similarity and semantic consistency. Additional results of ablations on individual layout control and textual–visual components are reported in Tabs 3 and 4. More quantitative and qualitative result with further analyses will be provided in Appendix.
@@ 4.4 Downstream Application
We further evaluate the impact of synthesized camouflaged images on downstream COD application. An awesome dichotomous segmentation model, MVANet (CVPR2024MVANet), is trained with different data configurations and evaluated on four COD benchmarks. Quantitative results and prediction results are presented in Tab. 5 and Fig. 6. We can observe that, converting salient or general images into camouflaged ones degrades COD performance, possibly due to domain gaps between these synthesized samples and real COD data. In contrast, training with camouflaged-synthesis images consistently improves COD performance, highlighting the effectiveness of our generation approach for data augmentation. We also examine the impact of synthesized images on salient and general object detection. More details of experimental settings, metrics and results, and the constructed SynCOD12K dataset, will be reported in Appendix.
@@ 5 Conclusion
In this paper, we propose RealCamo, a unified framework for controllable camouflaged image generation. RealCamo explicitly incorporates additional layout controls to regulate the structural composition of generated images, thereby enhancing semantic consistency between foreground objects and synthesized backgrounds. In addition, it constructs a multi-modal textual–visual condition by combining unified fine-grained textual task descriptions with texture-oriented background retrieved visual cues , which jointly guides the generation process to improve visual quality and realism. Extensive experiments and qualitative analyses validate the effectiveness of the proposed framework and demonstrate its strong potential for enhancing downstream camouflaged object detection applications.
Appendix A Camouflage Object Detection

### v2

@@ 3.4 Texture-Oriented Background Retrieval
Compared to generic image synthesis, camouflaged image generation requires backgrounds that are quite complex and sometimes unconventional. Although LAKE-RED (Zhao et al., 2024) retrieves background embeddings from a pretrained codebook, such representations learned from generic image distributions lack camouflage-specific priors. We therefore introduce a Texture-Oriented Background Retrieval (TOBR) module that uses camouflaged images as a domain-specific background knowledge base. As illustrated in Algorithm 1, given a target foreground object, a CLIP image encoder is adopted to extract texture-aware features for the object and all candidates in the knowledge base. Then, the top-$k$ most similar images in the embedding space are retrieved and incorporated into multimodal prompt construction to guide the camouflage generation.
@@ 3.5 Textual-Visual Condition Generator
Textual conditioning is central to T2I generation. Although Das and Gopalakrishnan (2025) enhances textual prompts via prompt engineering, the incorporated background semantics remain limited. While detailed scene descriptions have been shown to improve the realism of camouflaged image synthesis (Qian et al., 2025; Chen et al., 2026), acquiring scene-specific fine-grained descriptions is often labor-intensive. We therefore propose a Textual–Visual Condition Generator (TVCG) that constructs a multimodal prompt by combining a unified fine-grained task description with high-quality background visual embeddings. This design offers richer guidance for generation and improves visual similarity between the foreground target and the background. The multimodal prompt comprises three components: (i) fine-grained textual task description, (ii) object category information, and (iii) reference background visual cues.
Fine-Grained Textual Task Description. Instead of using overly simplistic prompts (e.g., “image of a fish Camouflaged with Background” (Das and Gopalakrishnan, 2025)) or labor-intensive, scene-specific descriptions (e.g., “A fish with mottled green and brown patterns and elongated appendages in a coral reef ecosystem.” (Qian et al., 2025)), we adopt a unified fine-grained textual task description “A realistic image of an object blending into its surroundings, where the background shares similar colors, textures, and patterns with the object, making it hard to distinguish. Natural lighting, photorealistic, seamless camouflage, high detail.” as the textual guidance in generation. This unified description improves generation quality while eliminating the need for manually crafting scene-specific prompts, and is encoded into textual tokens $c_{\mathrm{txt}}$ via a CLIP text encoder.
Object Category Information. For the given image, we adopt the [CLS] token processed by a CLIP image encoder as a compact category representation $c_{\mathrm{cls}}$ of the foreground object, avoiding explicit class identification while preserving informative semantic guidance.
@@
Reference Background Visual Cues. To provide richer background guidance for the camouflage generation process, we incorporate the samples retrieved by TOBR as visual reference knowledge into the multimodal prompt. Specifically, we employ a CLIP image encoder to extract visual embeddings from the input image $\mathrm{I}$ and its top-$k$ TOBR-retrieved samples $\{S_{j}\}_{j=1}^{k}$. Guided by the foreground mask $\mathrm{M}$, the foreground region of the input image is preserved, while the background regions are fused with the retrieved visual knowledge to construct background-aware visual cues $c_{\mathrm{vis}}$. This process can be formulated as:
@@
$c_{\mathrm{vis}}=\left\{\begin{aligned} &e_{t}\cdot\mathrm{M}+\frac{(\mathbf{1}-\mathrm{M})}{k+1}\cdot(e_{t}+\sum_{j=1}^{k}e_{j}),\ \textbf{if}\ \ training\\ &e_{t}\cdot\mathrm{M}+\frac{(\mathbf{1}-\mathrm{M})}{k}\cdot\sum_{j=1}^{k}e_{j},\ \textbf{if}\ \ inference.\end{aligned}\right$. (5)
The final multimodal conditional prompt is the concatenation of $_{\mathrm{txt}}, cc$$_{\mathrm{cls}}$, and $c_{\mathrm{vis}}$ along the sequence dimension.
Figure 4: Qualitative comparison with existing SOTA methods including style transfer-based (column 3 to 4), out-painting-based (column 5 to 9), and text-driven (column 10) approaches. Our results are demonstrated in the rightmost column with red box.
Table 1: Quantitative comparison with SOTA methods on LAKE-RED datasets. The 1st, 2nd and 3rd best results are highlighted in red, green and blue. $\uparrow$ / $\downarrow$ represents the higher / lower the better. ST: Style transfer-based. OP: Out-paining-based. TD: Text-driven.
@@ 3.6 Training Objective
In RealCamo, the Stable Diffusion (SD) and the CLIP text and image encoders are all frozen, leaving ControlNet as the only trainable component. The training objective follows Equation 2, where the original control $c^{f}$ is replaced by layout controls $\{c_{\mathrm{cst}},c_{\mathrm{dep}},c_{\mathrm{hed}}\}$, and the textual condition $c^{t}$ is replaced by multimodal conditions $\{c_{\mathrm{txt}},c_{\mathrm{cls}},c_{\mathrm{vis}}\}$. This design enables effective controllable camouflage synthesis while preserving the generalization of the pretrained SD.
@@ 3.7 Background-Foreground Distribution Divergence
To more effectively quantify the degree of camouflage in synthesized images, we propose a background–foreground distribution divergence metric, denoted as $\mathrm{KL}_{BF}$, based on KL divergence. It measures visual similarity between the generated background and the foreground targets via their pixel-value distributions, where lower values indicate stronger camouflage effectiveness. Specifically, given an image $\mathrm{I}$ and its foreground mask $\mathrm{M}$, we compute the KL divergence between background and foreground distributions for each RGB channel and average them to obtain $\mathrm{KL}_{BF}$. Formally, the $\mathrm{KL}_{BF}$ can be calculated as:
$\mathrm{KL}_{c}$ $=KL(\mathrm{I}_{c}\cdot(\mathbf{1}-\mathrm{M}),\mathrm{I}_{c}\cdot\mathrm{M}),$ (6)
$\mathrm{KL}_{BF}$ $=Avg(\{\mathrm{KL}_{c}\}),\ c\in\{R,G,B\}.$ (7)
Figure 5: Qualitative results of the ablation study on main components of our proposed RealCamo framework.
Table 2: Quantitative results of the ablation study on main components. The best results are highlighted in bold. $\uparrow$ / $\downarrow$ represents the higher / lower the better. For layout controls, $\times$ means only a foreground target with a white background is used, without depth and HED controls. For the textual-visual condition, $\times$ means only a simple textual prompt is used, without additional visual tokens.
@@ 4.1 Experimental Settings
Datasets and Metrics. We conduct experiments on the LAKE-RED dataset (Zhao et al., 2024), which contains 4,040 training images and 19,419 images for evaluation, including 6,473 camouflaged, salient, and general images, respectively. Following Zhao et al. (2024), we adopt FID (Binkowski et al., 2018) and KID (Heusel et al., 2017) to assess the overall quality of generated images. In addition, SSIM (Wang et al., 2004) is employed to measure structural consistency between generated images and their original references, while our proposed $\mathrm{KL}_{BF}$ metric is used to quantify the effectiveness of visual camouflage.
Implementation Details. The proposed framework is built upon Stable Diffusion 1.5 (Rombach et al., 2022) augmented with ControlNet (Agrawala, 2023). During training, all input images, layout controls, and corresponding masks are resized to $512\times 512$, and a batch size of 4 is used. We set the control scale to 1.0 and train the model for 600,000 iterations with a learning rate of 1e-5. The generated images are also in a resolution of $512\times 512$ and are resized to their original size for evaluation. All experiments are conducted on a single NVIDIA L20 GPU.
@@ 4.2 Comparison with the State-of-the-arts (SOTAs)
We compare our RealCamo with 12 SOTAs, including five style transfer-based methods (AB (Martino et al., 2016), CI (Chu et al., 2010), AdaIN (Huang and Belongie, 2017), DCI (Zhang et al., 2020), and LCGNet (Li et al., 2022b)), six out-painting-based methods (TFill (Zheng et al., 2022), LDM (Rombach et al., 2022), RePaint-L (Lugmayr et al., 2022), LAKE-RED (Zhao et al., 2024), FACIG (Chen et al., 2025), and CamoAny (Das and Gopalakrishnan, 2025)), and one recently text-driven method (CT-CIG (Qian et al., 2025)). We re-evaluate the benchmark results provided by LAKE-RED and CT-CIG. Due to a lack of available resources, we reproduce CamoAny with the same settings as ours, and directly copy the results from the original paper for FACIG.
Table 5: Performance of MVANet (Yu et al., 2024a) on four COD benchmark datasets with different training data. The best results are highlighted in bold. $\uparrow$ / $\downarrow$ represents the higher / lower the better. Original (Ori.): The standard COD training data. Gen S&G: Synthesis images by RealCamo of the salient and general subsets in the LAKE-RED validation set. SynCOD12K: Synthesis images by RealCamo of the LAKE-RED training set.
@@
Ori. + Gen S&G 0.905 0.927 0.846 0.796 0.031 0.826 0.842 0.760 0.717 0.079 0.865 0.896 0.784 0.719 0.035 0.879 0.897 0.821 0.769 0.048
Ori. + SynCOD12K 0.924 0.949 0.889 0.871 0.019 0.882 0.916 0.856 0.826 0.049 0.896 0.934 0.850 0.818 0.021 0.891 0.919 0.860 0.824 0.037
Qualitative Comparison. Fig. 4 shows the qualitative comparison between our RealCamo and other methods on camouflaged, salient, and generic objects. The style transfer-based LCGNet embeds targets into given backgrounds, achieving strong visual camouflage but severely altering target semantics. In contrast, out-painting-based methods often produce insufficient camouflage or cluttered, semantically implausible backgrounds, degrading realism. Although recent text-driven CT-CIG generates visually convincing results, they fail to preserve original targets and thus require re-annotation, limiting their usefulness for downstream COD tasks. In comparison, our RealCamo achieves both high visual similarity and strong foreground–background semantic consistency while preserving target integrity, resulting in more realistic and practically valuable camouflaged images.
Quantitative Comparison. Quantitative results are summarized in Tab. 1. Compared to early style transfer-based methods, recent out-painting-based and text-driven approaches achieve obviously better image quality, indicated by lower FID and KID. Owing to the introduction of explicit layout controls and textual–visual multimodal conditioning, our RealCamo attains highly competitive generation performance. Although it does not achieve the lowest $\mathrm{KL}_{BF}$ score, our method balances visual camouflage with structural and semantic fidelity, producing more realistic images. By revisiting Fig. 4, methods with lower $\mathrm{KL}_{BF}$ values (e.g., CamoAny and TFill) indeed yield stronger visual camouflage, further validating our proposed $\mathrm{KL}_{BF}$ as an effective metric for quantifying camouflage effectiveness.
User Study. We further conduct a user study to comprehensively evaluate the quality of synthesized camouflaged images based on human feedback. Detailed experimental settings, evaluation criteria, and statistical results and analyses are reported in Appendix E.
@@ 4.3 Ablation Study
We conduct ablation studies to validate the proposed components. Due to computational constraints, all variants are trained for 100,000 iterations and evaluated on 6,000 images uniformly sampled from three validation subsets. Results in Tab. 2 and Fig. 5 show that introducing explicit layout controls improves structural consistency and overall image quality, supporting our claim that “the semantics of images are largely determined by their structural layout”. Incorporating textual–visual condition further enhances visual realism and foreground–background visual similarity and semantic consistency. Additional results of ablations on individual layout control and textual–visual components are reported in Tabs 3 and 4. More quantitative and qualitative results with further analyses are provided in Appendix C.
@@ 4.4 Downstream Application
We further evaluate the impact of synthesized camouflaged images on the downstream COD task. An awesome dichotomous segmentation model, MVANet (Yu et al., 2024a), is trained with different data configurations and evaluated on four COD benchmarks. Quantitative results and prediction results are presented in Tab. 5 and Fig. 6. We can observe that converting salient or general images into camouflaged ones degrades COD performance, possibly due to domain gaps between these synthesized samples and real COD data. In contrast, training with camouflaged-synthesis images consistently improves COD performance, highlighting the effectiveness of our generation approach for data augmentation. We also examine the impact of synthesized images on salient and general object detection. More details of experimental settings, metrics and results, and the constructed SynCOD12K dataset, are reported in Appendix D.
@@ 5 Conclusion
In this paper, we propose RealCamo, a novel framework for controllable realistic camouflaged image generation. RealCamo explicitly incorporates additional layout controls to regulate the structural composition of generated images, thereby enhancing semantic consistency between foreground objects and synthesized backgrounds. In addition, it constructs a multimodal textual–visual condition by combining a unified fine-grained textual task description with texture-oriented background retrieved visual cues, which jointly guide the generation process to improve visual quality and realism. Extensive experiments and qualitative analyses validate the effectiveness of the proposed framework and demonstrate its strong potential for enhancing downstream camouflaged object detection applications.
Appendix A Camouflaged Object Detection